OpenStudy (tanabugg):

write a function rule to represent the sequence 47,32,17,2 a. A(n)= 47 +(n - 1)15 b. A(n)= 47 +(n - 1)(-13) c. A(n)= 47+(n - 1)(-15) d. A(n)= 32 + (n - 1)15

OpenStudy (nintendo):

Just substitute values of n into all of these functions. In this case, n=2 is sufficient to give you a correct answer. For option a, A(2)=47+15=62, while 32 is expected, so it is incorrect. For option b, A(2)=47-13=35, so it is incorrect. For option d, A(n)=47, so it is incorrect. Option c, however, gives A(2)=47-15=32, so it is valid. As we have proven that all of the other options are not valid, (c) is correct.

OpenStudy (tanabugg):

thanks!!!!!